OpenStudy (anonymous):

(5x^4 +6)^2 how do I square the expression?

Parth (parthkohli):

I think you must use the identity right here. \(\Rightarrow (5x^4 + 6)^2 = (5x^4)^2 + 2(6)(5x^4) + 6^2\)

OpenStudy (anonymous):

(a+b)^2- a^2 + 2ab+ b^2 like that??

OpenStudy (anonymous):

\[(5^{4} + 6){2} = (5x{4} + 6)(5x{4} + 6)\]

OpenStudy (anonymous):

its an identity lol

OpenStudy (radar):

@jk_16, what did you get for your final answer.??

OpenStudy (radar):

The question could not be answered by anyone else but @jk_16, because it was asked how you would square the expression lol

OpenStudy (radar):

The way I would square it is probaby different from the way you would. Excluding the fact it is a "Special Product" I would use the old fashion way.

OpenStudy (anonymous):

a=5x^4 b=6 (5x^4)^2 + 2*6*5x^4 + 6^2 how I did it

OpenStudy (radar):